# CliffordNet: All You Need is Geometric Algebra

### v1

@@ Abstract
Implemented via an efficient sparse rolling mechanism with strict linear complexity $\mathcal{O}(N)$, our model reveals a surprising emergent property: the geometric interaction is so representationally dense that standard Feed-Forward Networks (FFNs) become redundant. Empirically, CliffordNet establishes a new Pareto frontier: our Nano variant achieves 76.41% accuracy on CIFAR-100 with only 1.4M parameters, effectively matching the heavy-weight ResNet-18 (11.2M) with $8\times$ fewer parameters, while our Base variant sets a new SOTA for tiny models at 78.05%. Our results suggest that global understanding can emerge solely from rigorous, algebraically complete local interactions, potentially signaling a shift where geometry is all you need. Code is available at https://github.com/ParaMind2025/CAN.
@@ 3.2 A General Framework for Feature Evolution
We model the layer-wise update of visual features as a continuous dynamic evolution process. Specifically, we hypothesize that the evolution of $H$ follows a differential equation governed by the interaction between the current state and its local geometric context:
@@ 3.3 The Clifford Interaction Ansatz
3.4 Context Instantiation via Discrete Laplacian
While $\mathcal{C}(H)$ could theoretically take many forms, we posit that the most fundamental geometric context for visual data is local variation (e.g., edges and texture gradients). In vector calculus, this is naturally captured by the Laplacian operator $\Delta$.
The Generalized Discrete Laplacian. In the discrete Euclidean grid of an image, the Laplacian $\Delta H$ is typically computed via finite difference filters (e.g., the 5-point stencil). Instead of hand-crafting these filters, we employ depth-wise convolutions to learn a diverse basis of generalized difference operators for each channel. This allows the network to adaptively capture spatial derivatives along different orientations and frequencies, providing a rich description of the local geometry.
Iterative Diffusion via Factorization. To capture macroscopic geometric structures and texture gradients rather than just microscopic noise, a large receptive field is essential, aligning with recent findings in modern ConvNet design (e.g., $7\times 7$ kernels) [10]. However, rather than employing a single large kernel, we find that a factorized approach yields superior results. We define the context stream as a stack of smaller depth-wise convolutions (e.g., $3\times 3$):
$\mathcal{C}(H)\coloneqq\Delta_{\text{iter}}H$ (3)
where the operator $\Delta_{\text{iter}}$ denotes an iterative finite difference scheme. We instantiate this operator as a composition of local difference filters (implemented via stacked $3\times 3$ depth-wise convolutions). Mathematically, this factorization $\Delta_{\text{iter}}\equiv\mathcal{K}_{2}*(\mathcal{K}_{1}*\cdot)$ approximates a macroscopic Laplacian with an expanded effective receptive field ($5\times 5$), akin to performing multi-step numerical integration of a diffusion process.
Hardware-Aware Implementation. By approximating the manifold operator via stacked standard $3\times 3$ convolutions, our design is hardware-aware. It leverages the highly optimized convolution primitives in modern accelerators (e.g., cuDNN on GPUs), ensuring high throughput and memory efficiency, in contrast to theoretically precise but computationally fragmented operators like sparse graph convolutions.
Under this instantiation, Eq. (1) effectively describes a geometric evolution system where features evolve not just by smoothing, but by interacting geometrically with their own iteratively refined local curvature:
@@ 3.4 Context Instantiation via Discrete Laplacian
where the operators $\mathcal{D}(\cdot,\cdot)$ and $\mathcal{W}(\cdot,\cdot)$ correspond to the Generalized Inner Product and Exterior Product components of the Clifford geometric product, respectively. Physically, $\mathcal{D}$ acts as a coherence force (aligning features), while $\mathcal{W}$ acts as a structural torque (capturing orthogonality).
@@ 3.5 Efficient Realization: The Shifted Geometric Product
Interpretation. The Dot term (Eq. 7 top) corresponds to a generalized inner product $\mathbf{u}^{\top}\mathbf{P}_{s}\mathbf{v}$, where $\mathbf{P}_{s}$ is the cyclic permutation matrix corresponding to shift $s$. Enhanced with the SiLU activation, it acts as a non-linear coherence gate, filtering features based on their alignment with the shifted context. The Wedge term (Eq. 7 bottom) explicitly computes the Bivector Coefficients for the plane spanned by basis vectors $\mathbf{e}_{c}$ and $\mathbf{e}_{c+s}$. Geometrically, this term represents the oriented area of the parallelogram spanned by the feature vectors. This area is maximized when features are orthogonal, allowing the network to explicitly detect structural variations and edges. The Wedge term explicitly computes the bivector coefficients ($h_{i}c_{i+s}-c_{i}h_{i+s}$), strictly adhering to the anti-symmetric definition of the exterior product to capture geometric orientation.
@@
Mathematically, the shifted dot product term $\mathbf{u}\cdot(\mathcal{T}_{s}\mathbf{v})$ can be formalized as a bilinear form $\mathbf{u}^{\top}\mathbf{P}_{s}\mathbf{v}$. Unlike the standard inner product (where $\mathbf{P}_{s}=\mathbf{I}$), this formulation generalizes the notion of similarity. Here, $\mathbf{P}_{s}$ acts as an adjacency operator defining a specific topological connectivity between channels. This allows the network to measure coherence not just between identical features, but across semantically related yet spatially distant feature channels encoded at specific intervals $s$.
@@ 3.6 Gated Geometric Residual (GGR)
3.7 Architectural Variants
To investigate the limits of geometric efficiency, we instantiate CliffordNet in two configurations:
  1. 1.
    CliffordNet-Base: The standard configuration equipped with a lightweight FFN (MLP ratio = 1.0) after the geometric block, providing a balance between geometric interaction and explicit channel mixing.
  2. 2.
    CliffordNet-Fast (No-FFN): An extreme efficiency variant where the FFN is completely removed ($\text{mlp\_ratio}=0$). In this setting, the network relies solely on the Clifford Interaction and GGR fusion for all feature transformations. The complete forward pass of this variant is detailed in Algorithm 1.
We provide a theoretical analysis of why the No-FFN design remains effective in Section 5.
Algorithmic Summary. To provide a holistic view of this paradigm shift, we present the complete forward pass of the CliffordNet Block in Algorithm 1. This algorithm explicitly details the No-FFN variant, illustrating how the distinct geometric components — Dual-Stream generation, Sparse Rolling Interaction, and Gated Geometric Residual — orchestrate to evolve features. Notably, the workflow terminates immediately after the geometric fusion (Line 34), bypassing the computational and parameter overhead of the traditional MLP block used in MetaFormer architectures.
Algorithm 1 Pseudo-code for CliffordNet Block (No-FFN Variant)
@@ 3.7 Architectural Variants
6:  /* 3. Efficient Interaction (Rolling) */
7:  $\mathbf{F}_{list}\leftarrow[]$
8:  for $s$ in $\mathcal{S}$ (e.g., {1, 2, 4, …}) do
9:   $\mathbf{Z}_{d}^{(s)}\leftarrow\mathbf{Roll}(\mathbf{Z}_{det},s)$
10:   $\mathbf{Z}_{c}^{(s)}\leftarrow\mathbf{Roll}(\mathbf{Z}_{ctx},s)$
11:   /* Geometric Product */
12:   $\mathbf{W}^{(s)}\leftarrow\mathbf{Z}_{det}\odot\mathbf{Z}_{c}^{(s)}-\mathbf{Z}_{ctx}\odot\mathbf{Z}_{d}^{(s)}$
13:   $\mathbf{D}^{(s)}\leftarrow\mathbf{SiLU}(\mathbf{Z}_{det}\odot\mathbf{Z}_{c}^{(s)})$
14:   $\mathbf{F}_{list}.\mathbf{append}(\mathbf{W}^{(s)},\mathbf{D}^{(s)})$
15:  end for
15:  /* Concat & Project back to $\mathbb{R}^{C}$ */
16:  $\mathbf{G}_{raw}$ : $(\dots,\mathtt{2|\mathcal{S}|C})$ $\leftarrow$ $\mathbf{Cat}(\mathbf{F}_{list})$
17:  $\mathbf{G}_{feat}$ : $(\dots,\mathtt{C})$ $\leftarrow$ $\mathbf{Linear}_{proj}(\mathbf{G}_{raw})$
18:  /* 4. Gated Geometric Residual */
19:  $\mathbf{M}$ $\leftarrow$ $\mathbf{Cat}([\mathbf{X}_{ln},\mathbf{G}_{feat}])$
20:  $\mathbf{\alpha}$ $\leftarrow$ $\mathbf{Sigmoid}(\mathbf{Linear}_{gate}(\mathbf{M}))$
21:  /* Non-linear modulation */
22:  $\mathbf{H}_{mix}$ $\leftarrow$ $\mathbf{SiLU}(\mathbf{X}_{ln})+\mathbf{\alpha}\odot\mathbf{G}_{feat}$
23:  /* 5. Output Update */
24:  /* Note: No FFN applied */
25:  $\mathbf{X}_{l}$ $\leftarrow$ $\mathbf{X}_{l-1}+\mathbf{Drop}(\mathbf{\gamma}\odot\mathbf{H}_{mix})$
26:  Return: $\mathbf{X}_{l}$
27:
@@ 4 Experiments
Table 1: Main Results on CIFAR-100. We compare CliffordNet against established efficient backbones under an identical training recipe. Note on Speed: Baseline CNNs benefit from mature cuDNN kernel optimizations. Despite relying on unoptimized PyTorch primitives (e.g., rolling), CliffordNet-Nano remains highly competitive in training time while delivering significantly superior accuracy (+1.8% vs. ShuffleNet, +5.5% vs. MobileNetV2) with fewer parameters.
@@
CliffordNet-Fast 2.6M No 5 125.0 77.63%
CliffordNet-Base 3.0M Yes 5 139.0 78.05%

### v2

@@ Abstract
Implemented via an efficient sparse rolling mechanism with strict linear complexity $\mathcal{O}(N)$, our model reveals a surprising emergent property: the geometric interaction is so representationally dense that standard Feed-Forward Networks (FFNs) become redundant. Empirically, CliffordNet establishes a new Pareto frontier: our Nano variant achieves 77.82% accuracy on CIFAR-100 with only 1.4M parameters, effectively matching the heavy-weight ResNet-18 (11.2M) with $8\times$ fewer parameters, while our Lite variant (2.6M) sets a new SOTA for tiny models at 79.05%. Our results suggest that global understanding can emerge solely from rigorous, algebraically complete local interactions, potentially signaling a shift where geometry is all you need. Code is available at https://github.com/ParaMind2025/CAN.
@@ 3.2 A General Framework for Feature Evolution
We model the layer-wise update of visual features as a continuous dynamic evolution process. Specifically, we hypothesize that the evolution of $H$ follows a differential equation governed by the interaction between the current state and its geometric context:
@@ 3.3 The Clifford Interaction Ansatz
Primer on Bivector Bases. To demystify the exterior product, consider the input features in a standard basis $\{\mathbf{e}_{1},\dots,\mathbf{e}_{D}\}$. The wedge product $\mathbf{u}\wedge\mathbf{v}$ yields a bivector, which can be expanded as a linear combination of basis bivectors:
$\mathbf{u}\wedge\mathbf{v}=\sum_{1\leq i<j\leq D}(u_{i}v_{j}-v_{i}u_{j})(\mathbf{e}_{i}\wedge\mathbf{e}_{j})$ (3)
Here, each basis element $\mathbf{e}_{i}\wedge\mathbf{e}_{j}$ represents a unit oriented plane spanning dimensions $i$ and $j$. While the full bivector space has dimension $\binom{D}{2}$, our architecture does not compute all coefficients. Instead, it selectively computes those corresponding to specific spatial/channel relationships defined by the rolling shifts, ensuring efficiency while capturing essential geometric structures.
@@
3.4 Context Instantiation: Local-Global Duality
The geometric context $\mathcal{C}(H)$ serves as the reference field driving feature evolution. In our framework, $\mathcal{C}$ is not constrained to a single scale; rather, it can be instantiated to capture either local structural variation or global semantic coherence, depending on the design constraints.
  • Local Context (Implicit Globality): For extreme efficiency, $\mathcal{C}$ can be instantiated via local operators (e.g., convolutions). Here, the global receptive field emerges implicitly through the layer-wise propagation of local interactions, akin to a diffusion process.
  • Global Context (Explicit Globality): To maximize performance, $\mathcal{C}$ can be derived from a global summary (e.g., global average, linear attention). This mirrors the global information flow in Transformers but differs fundamentally in complexity. While Transformers rely on quadratic Attention ($\mathcal{O}(N^{2})$), our Global Clifford Interaction remains strictly linear ($\mathcal{O}(N)$), interacting the local state $H_{i}$ with the global mean field $\mathcal{C}_{glo}$.
Design Choice. We posit that the Global Context variant is the natural geometric successor to the standard Feed-Forward Network (FFN). While FFNs mix channels blindly, the Global Clifford Interaction aligns local features with the global scene context geometrically. Thus, for our high-performance variants, we employ a superposition of both scales.
1. Instantiating the Local Context (Diffusion).
Factorized Linear Laplacian (The Smoothing Step). To capture macroscopic structures with a large receptive field (e.g., $7\times 7$) while optimizing parameter efficiency, we adopt a factorized strategy. We define the local context field $\mathcal{C}_{loc}$ as a stack of two smaller depth-wise convolutions separated by non-linear activation:
$\mathcal{C}_{loc}(H)=\text{Conv}_{3\times 3}\Big(\text{Conv}_{3\times 3}(H)\Big)$ (4)
This design is hardware-aware: it leverages highly optimized $3\times 3$ convolution primitives (e.g., cuDNN) to ensure high throughput, avoiding the latency overhead of sparse or irregular operators.
Generalized Geometric Prior (The Adaptation Step). We further generalize the context definition by introducing a self-energy suppression factor $\lambda\in\{0,1\}$. The final context $\mathcal{C}$ entering the interaction layer is defined as:
$\mathcal{C}=\mathcal{C}_{loc}(H)-\lambda\cdot H$ (5)
This formulation unifies two distinct geometric modes based on model capacity:
  • Differential Mode ($\lambda=1$): Here, the context effectively becomes the discrete Laplacian $\mathcal{C}\approx\Delta H$. This formulation acts as a pure geometric high-pass filter, explicitly suppressing the base feature intensity from the identity path. We find this optimal for capacity-constrained models, as it maximizes the signal-to-noise ratio of structural variations.
  • Absolute Mode ($\lambda=0$): $\mathcal{C}=\mathcal{C}_{loc}$. This acts as an energy-preserving flow. The subsequent inner product ($H\cdot\mathcal{C}\approx\|H\|^{2}+H\cdot\Delta H$) retains feature intensity information.
2. Global-Local Geometric Superposition
While the local context effectively captures intricate local geometric structures, integrating the global context can further enhance the overall representation. To achieve this without the quadratic cost of self-attention, we introduce a Dual-Scale Superposition mechanism.
We define a global context stream $\mathcal{C}_{glo}$ derived from the global average of the feature field:
$\mathcal{C}_{glo}=\text{GlobalAvgPool}(H)$ (6)
We then compute the Clifford interaction between the state $H$ and this global context. This operation, which we term geometric FFN-Global (gFFN-G), serves as a geometrically interpretable alternative to standard MLPs. gFFN-G computes the full geometric product $H\mathcal{C}_{glo}$. The Wedge term ($H\wedge\mathcal{C}_{glo}$) specifically highlights features that deviate from the global background (saliency detection), while the Inner term ($H\cdot\mathcal{C}_{glo}$) reinforces features consistent with the global scene.
Unified Superposition Principle. Following the physical principle of field superposition, we integrate the local and global geometric forces via a weighted summation. The governing evolution equation is thus generalized as:
$\frac{\partial H}{\partial t}=\mathcal{P}_{loc}\Big(H(\mathcal{C}_{loc})\Big)+\beta\cdot\mathcal{P}_{glo}\Big(H(\mathcal{C}_{glo})\Big)$ (7)
where: $\mathcal{C}_{loc}\approx\Delta H$ represents the local differential context (via factorized convolutions). $\mathcal{C}_{glo}=\text{GlobalAvg}(H)$ represents the global mean field. $\beta\in\{0,1\}$ serves as a structural switch: $\beta=0$ prioritizes extreme parameter efficiency by relying solely on local geometric diffusion; $\beta=1$ integrates the global interaction mechanism to capture long-range semantic dependencies.
This design allows the network to simultaneously sharpen local boundaries (via local curvature) and highlight salient objects (via global contrast), unifying micro and macro perception in a single, linear-complexity update step.
3. Unified Algebraic Mechanism.
Crucially, despite the distinct physical interpretations of $\mathcal{C}_{loc}$ (local curvature) and $\mathcal{C}_{glo}$ (global mean field), both terms in Eq. (7) are governed by an identical algebraic mechanism: the Clifford Geometric Product.
Therefore, in the following sections, we focus our derivation on the general interaction form $\mathcal{F}(H,\mathcal{C})=\mathcal{P}(H(\mathcal{C}))$. Whether $H$ interacts with its neighbor or the global context, the underlying operation remains the same: a unification of the generalized inner product (coherence) and the exterior product (structure):
@@ 2. Global-Local Geometric Superposition
where the operators $\mathcal{D}(\cdot,\cdot)$ and $\mathcal{W}(\cdot,\cdot)$ correspond to the Generalized Inner Product and Exterior Product components of the Clifford geometric product, respectively.
@@ 3.5 Efficient Realization: The Shifted Geometric Product
Interpretation. The Inner term (Eq. 11 top) corresponds to a generalized inner product $\mathbf{u}^{\top}\mathbf{P}_{s}\mathbf{v}$, where $\mathbf{P}_{s}$ is the cyclic permutation matrix corresponding to shift $s$. Enhanced with the SiLU activation, it acts as a non-linear coherence gate, filtering features based on their alignment with the shifted context. The Wedge term (Eq. 11 bottom) explicitly computes the Bivector Coefficients for the plane spanned by basis vectors $\mathbf{e}_{c}$ and $\mathbf{e}_{c+s}$. Geometrically, this term represents the oriented area of the parallelogram spanned by the feature vectors. This area is maximized when features are orthogonal, allowing the network to explicitly detect structural variations and edges. The Wedge term explicitly computes the bivector coefficients ($h_{i}c_{i+s}-c_{i}h_{i+s}$), strictly adhering to the anti-symmetric definition of the exterior product to capture geometric orientation.
@@
Mathematically, the shifted inner product term $\mathbf{u}\cdot(\mathcal{T}_{s}\mathbf{v})$ can be formalized as a bilinear form $\mathbf{u}^{\top}\mathbf{P}_{s}\mathbf{v}$. Unlike the standard inner product (where $\mathbf{P}_{s}=\mathbf{I}$), this formulation generalizes the notion of similarity. Here, $\mathbf{P}_{s}$ acts as an adjacency operator defining a specific topological connectivity between channels. This allows the network to measure coherence not just between identical features, but across semantically related yet spatially distant feature channels encoded at specific intervals $s$.
@@
Matrix Interpretation: Circulant Sparsity. The rolling interaction can be rigorously interpreted through the lens of structured matrices. Conceptually, the full channel-wise interaction between state $H$ and context $\mathcal{C}$ forms a dense correlation matrix $\mathbf{M}\in\mathbb{R}^{D\times D}$, where $M_{i,j}$ represents the interaction between channel $i$ and $j$.
Our cyclic shift operation ($\mathcal{T}_{s}$) effectively samples specific diagonals from this matrix. If we considered all possible shifts $s\in\{0,\dots,D-1\}$, the operation would be equivalent to a multiplication by a Circulant Matrix (or Toeplitz matrix without wrapping). By restricting $s$ to a sparse set $\mathcal{S}$ (e.g., logarithmic steps), we enforce a Structured Circulant Sparsity. This implies that we approximate the full dense geometry by only retaining the most significant frequency bands in the channel interaction topology, reducing complexity from $\mathcal{O}(D^{2})$ to $\mathcal{O}(|\mathcal{S}|D)$.
@@ 3.6 Gated Geometric Residual (GGR)
3.7 The gFFN Family: A Generalized Interaction Module
Standard architectures typically decouple feature processing into spatial mixing (e.g., Attention, Conv) and channel mixing (FFN/MLP). We challenge this dichotomy by proposing the geometric Feed-Forward Network (gFFN).
gFFN acts as a Spatial-Channel Entangler. By leveraging the Clifford geometric product, it simultaneously performs channel mixing (via projection) and spatial modulation (via context interaction). We unify our architectural variants under this framework by defining three modes of geometric context $\mathcal{C}$:
  • gFFN-L (Local): $\mathcal{C}=\Delta H$. Utilizes local difference operators to capture high-frequency texture and edges. This powers our Nano variant, maximizing efficiency.
  • gFFN-G (Global): $\mathcal{C}=\text{GlobalAvg}(H)$. Utilizes the global mean field to enforce semantic coherence.
  • gFFN-H (Hybrid): $\mathcal{C}=\Delta H+\beta\cdot\text{GlobalAvg}(H)$. Combines micro and macro contexts via field superposition. This powers our CliffordNet variants, achieving SOTA performance by integrating multi-scale geometric priors.
This modular design allows CliffordNet — and potentially other architectures — to flexibly trade off between local precision and global awareness based on computational constraints.
3.8 Algorithmic Summary
To provide a holistic view of this paradigm shift, we present the complete forward pass of the CliffordNet Block in Algorithm 1. This algorithm explicitly details the No-FFN variant, illustrating how the distinct geometric components — Dual-Stream generation, Sparse Rolling Interaction, and Gated Geometric Residual — orchestrate to evolve features. Notably, the workflow terminates immediately after the geometric fusion (Line 34), bypassing the computational and parameter overhead of the traditional MLP block used in MetaFormer architectures.
Algorithm 1 Pseudo-code for CliffordNet Block (No-gFFN-G Variant)
@@ 3.8 Algorithmic Summary
6:  if $\text{ctx\_mode}=^{\prime}\!\text{diff}\,^{\prime}$ then
7:   $\mathbf{Z}_{ctx}\leftarrow\mathbf{Z}_{ctx}-\mathbf{Z}_{det}$
8:  end if
9:  /* 3. Efficient Interaction (Rolling) */
10:  $\mathbf{F}_{list}\leftarrow[]$
11:  for $s$ in $\mathcal{S}$ (e.g., {1, 2, 4, …}) do
12:   $\mathbf{Z}_{d}^{(s)}\leftarrow\mathbf{Roll}(\mathbf{Z}_{det},s)$
13:   $\mathbf{Z}_{c}^{(s)}\leftarrow\mathbf{Roll}(\mathbf{Z}_{ctx},s)$
14:   $\mathbf{W}^{(s)}\leftarrow\mathbf{Z}_{det}\odot\mathbf{Z}_{c}^{(s)}-\mathbf{Z}_{ctx}\odot\mathbf{Z}_{d}^{(s)}$
15:   $\mathbf{D}^{(s)}\leftarrow\mathbf{SiLU}(\mathbf{Z}_{det}\odot\mathbf{Z}_{c}^{(s)})$
16:   /* Interaction Mode */
17:   if $\text{cli\_mode}=^{\prime}\!\text{inner}\,^{\prime}$ then
18:    $\mathbf{F}_{list}.\mathbf{append}(\mathbf{D}^{(s)})$
19:   else if $\text{cli\_mode}=^{\prime}\!\text{wedge}\,^{\prime}$ then
20:    $\mathbf{F}_{list}.\mathbf{append}(\mathbf{W}^{(s)})$
21:   else if $\text{cli\_mode}=^{\prime}\!\text{full}\,^{\prime}$ then
22:    $\mathbf{F}_{list}.\mathbf{append}(\mathbf{W}^{(s)},\mathbf{D}^{(s)})$
23:   end if
24:  end for
25:  /* Concat & Project back to $\mathbb{R}^{C}$ */
26:  if $\text{cli\_mode}=^{\prime}\!\text{full}\,^{\prime}$ then
27:   $\mathbf{G}_{raw}$ : $(\dots,\mathtt{2|\mathcal{S}|C})$ $\leftarrow$ $\mathbf{Cat}(\mathbf{F}_{list})$
28:  else if $\text{cli\_mode}\;\text{in}\;\{^{\prime}\!\text{inner}\,^{\prime},^{\prime}\!\text{wedge}\,^{\prime}\}$ then
29:   $\mathbf{G}_{raw}$ : $(\dots,\mathtt{|\mathcal{S}|C})$ $\leftarrow$ $\mathbf{Cat}(\mathbf{F}_{list})$
30:  end if
31:  $\mathbf{G}_{feat}$ : $(\dots,\mathtt{C})$ $\leftarrow$ $\mathbf{Linear}_{proj}(\mathbf{G}_{raw})$
32:  /* 4. Gated Geometric Residual */
33:  $\mathbf{M}$ $\leftarrow$ $\mathbf{Cat}([\mathbf{X}_{ln},\mathbf{G}_{feat}])$
34:  $\mathbf{\alpha}$ $\leftarrow$ $\mathbf{Sigmoid}(\mathbf{Linear}_{gate}(\mathbf{M}))$
35:  /* Non-linear modulation */
36:  $\mathbf{H}_{mix}$ $\leftarrow$ $\mathbf{SiLU}(\mathbf{X}_{ln})+\mathbf{\alpha}\odot\mathbf{G}_{feat}$
37:  /* 5. Output Update */
38:  /* Note: No FFN applied */
39:  $\mathbf{X}_{l}$ $\leftarrow$ $\mathbf{X}_{l-1}+\mathbf{Drop}(\mathbf{\gamma}\odot\mathbf{H}_{mix})$
40:  Return: $\mathbf{X}_{l}$
41:
@@ 4 Experiments
We employed a patch_size of 2 to achieve a favorable trade-off between efficiency and accuracy. While the results presented herein rely on the Differential Mode due to its demonstrated superiority on CIFAR-100, we acknowledge that different data distributions may benefit from distinct geometric priors. Therefore, we preserve both the Absolute and Differential modes in our framework, allowing for flexible adaptation to other datasets.
Table 1: Main Results on CIFAR-100. We compare CliffordNet against established efficient backbones under an identical training recipe. Note on Speed: Baseline CNNs benefit from mature cuDNN kernel optimizations. Despite relying on unoptimized PyTorch primitives (e.g., rolling), CliffordNet-Nano remains highly competitive in training time while delivering significantly superior accuracy with fewer parameters.
@@
CliffordNet-Lite 2.6M No 5 125.0 79.05%